| Operation Summary for March 19, 2012 - March 21, 2012 |
|
Wednesday March 21 The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI, MiniBooNE, SeaQuest, and FTBFT-1017 experiments taking beam. During the day shift, Operations turned off the Neutrino Area enclosures and SeaQuest at 8:33 AM in preparation for accesses. Operations halted beam to T-1017 at 8:44 AM so MI personnel could conduct slow spill studies to Switchyard (SY). Operations turned on the Muon Ring power supplies at 10:33 AM and Muon Ring personnel began a study. The NuMI beam permit tripped off at 10:50 AM when a power supply’s (V118) output drifted low; everything reset. The MI slow spill study ended at noon; Operations resumed sending beam to T-1017. SeaQuest ended their beamline work, came out of access, and resumed taking beam at 2:13 PM. During the evening shift, the Muon Ring study ended at 4:23 PM. T-1017 and SeaQuest made accesses during this shift. Thursday March 22 The midnight shift began with Operations sending beam to NuMI, MiniBooNE, SeaQuest, and T-1017. During the midnight shift, at 1:57 AM, the NuMI beam permit tripped off for the first of three times in about five hours due to V118; it reset. Operators adjusted the QXR at 2:57 AM. Operators added water to a Linac’s RF cavity water skid (LRF5) at 3:14 AM. A SY power supply (S:V100) tripped off at 7:22 AM; it had to be reset locally. T-1017 accessed during this shift. During the day shift, a Muon Ring person began a study at 8:41 AM. Operations prepared the Neutrino Area G2 enclosure and SeaQuest for accesses at 10:06 AM. Operations turned off NuMI at 10:10 AM in preparation for accessing MI-65 to look for a LCW leak. Controls personnel used this MI downtime to replace a MADC that might be the cause of the frequent V118 power supply trips. Operations reported that NuMI personnel found no leaks on the NuMI side of MI-65; Operators prepared MI for access at 12:36 PM. Operations prepared Booster for an access at 1 PM to let FESS personnel repair a sump pump float. Operations reported at 1:25 PM that MI and NuMI personnel found a LCW leak on Lambertson magnet in MI-62. MI personnel also found a leak from a MI-40 quadrupole magnet power supply, which a power supply tech will repair. Operations prepared SY enclosure C for an access at 1:39 PM. MI came out of access at 2:39 PM and Booster came out of access at 2:46 PM. Booster resumed sending beam around 3 PM. A Booster person began a study at 3:45 PM. During the evening shift, MI came back online and Operations established beam to NuMI at 4:38 PM and to T-1017 at 5:13 PM. Operations established beam to SeaQuest at 5:46 PM. The NuMI beam permit tripped off at 8:08 PM for the first of three times in two and half hours due to MIRF7 trips; everything reset. At 9:08 PM, the NuMI beam permit tripped off due to a quadrupole magnet power supply (E:QF107) drifting low; everything reset. At 9:52 PM, the NuMI beam permit tripped off for the first of two times in about two hours due to V118 trips; everything reset. Friday March 23 The midnight shift began with Operations sending beam to NuMI, MiniBooNE, SeaQuest, and T-1017. During the midnight shift, at 1:44 AM, the NuMI beam permit tripped off due to V118; everything reset. At 4:42 AM, a SY power supply (S:HT210) tripped off on a ground fault indication; Operators had to reset the power supply locally. Operations Coordinator (OpsCo) Plans for Today and the Weekend The OpsCo agrees with the Booster plans to access and replace a RF cavity water flow turbine and for Muon Ring studies for today. FTBF Meson experiment T-1012 will install as a parasitic experiment behind T-1017 this morning and begin taking beam later this afternoon; they will run all weekend. MTA personnel will also be taking beam later this afternoon. |
|
More Information Comments and Suggestions |
| last modified 03/23/2012 email Fermilab |
| Security, Privacy, Legal | Use of Cookies | |